Grade 11Mechanics

center of mass of two particles with masses 2kg and 1 kg located at(1,0,1)and (2,2,0) has the cordinates of.

bhargavi
10 Years agoGrade 11

1 Answer

Himanshu
10 Years ago
center of mass is m1.r1+m2.r2/m1+m2
hence putting values in the given questions center of the mass of the system
x coordinate: 2.1+1.2/2+1=4/3
y coordinate:2.0+1.2/2+1=2/3
zcoordinate: 2.1+1.0/2+1=2/3